OpenStudy (anonymous):

what is an equation for the line that passes through points (-1,-4) and (1,4). Write the equation in slope-intercept form.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-1\quad ,&-4)\quad &(1\quad ,&4) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\\\ \quad \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

I got 8=m(2)

OpenStudy (jdoe0001):

\(\bf slope = m= \cfrac{rise}{run} \implies \cfrac{4-(-4)}{1-(-1)} \implies \cfrac{8}{2}\implies 4\)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-1\quad ,&-4) \end{array} \\\quad \\ y-y_1=4(x-x_1)\) you could use either point really, I just happen to use the 1st one, the 2nd one will work just fine as well